OpenStudy (anonymous):

F x3 = (5-F) x7 where F is the force in Newtons. Determine the value of F ?

OpenStudy (gorv):

3f=35-7f 10f=35 f=3.5

OpenStudy (anonymous):

\[\large F\times 3=7(5-F)\] Expand the RHS [Right Hand Side] \[\large 3F=35-7F\] \[\large 10F=35\] \[\large F=\frac{7}{2}\]

OpenStudy (anonymous):

thanks for the prompt reply. took me a little longer than you guys but i got 3.5 too .cheers

OpenStudy (anonymous):

No worries mate.